### v3

# Inflations among quantum Grothendieck rings of type A

Abstract.
We introduce a collection of injective homomorphisms among the quantum Grothendieck rings of finite-dimensional modules over the quantum loop algebras of type $\mathrm{A}$. In the classical limit, it specializes to the inflation among the usual Grothendieck rings studied by Brito–Chari [J. Reine Angew. Math. 804, 2023]. We show that our homomorphisms respect the canonical bases formed by the simple $(q,t)$-characters, which in particular verifies a conjecture of Brito–Chari in loc. cit. We also discuss a categorification of our homomorphisms using the quiver Hecke algebras of type $\mathrm{A}_{\infty}$.
2020 Mathematics Subject Classification:
17B37, 81R50
1. Introduction
The monoidal category $\mathscr{C}$ of finite-dimensional representations of quantum loop algebra associated with a complex simple Lie algebra has been studied for over 30 years from various perspectives. As in the classical representation theory of simple Lie algebras, we have a highest weight type classification of simple representations in $\mathscr{C}$ due to Chari–Pressley [3]. However, beyond such a classification result, the situation is much more intricate. For instance, there is no known closed formula for the $q$-characters of general simple representations in $\mathscr{C}$.
The quantum Grothendieck ring $K_{t}(\mathscr{C})$, introduced by Nakajima [20] and Varagnolo–Vasserot [22] for type $\mathrm{ADE}$, and by Hernandez [8] for general type, is a one-parameter deformation of the Grothendieck ring $K(\mathscr{C})$. It carries a canonical basis formed by the $t$-analog of $q$-characters (or $(q,t)$-characters for short) of simple representations, constructed through the Kazhdan–Lusztig type algorithm. For type $\mathrm{ADE}$, Nakajima [20] applied his theory of quiver varieties to prove an analog of Kazhdan–Lusztig conjecture in this setting, that is, the simple $(q,t)$-characters specialize to the simple $q$-characters in the classical limit $t\to 1$. Thus, it gives an explicit algorithm to compute the simple $q$-characters of type $\mathrm{ADE}$ in a uniform way.
For the other type $\mathrm{BCFG}$, Hernandez, Oh, Oya, and the present author [6, 7] verified the analog of Kazhdan–Lusztig conjecture for some classes of simple representations, including all the simple representations of type $\mathrm{B}$. A crucial step in their proof is the construction of a collection of isomorphisms among the quantum Grothendieck rings of different Dynkin types, respecting the canonical bases. For example, it includes some isomorphisms between the quantum Grothendieck rings of type $\mathrm{B}_{n}$ and type $\mathrm{A}_{2n-1}$.
This result partly motivates us to investigate further relationship among the quantum Grothendieck rings. In this paper, we restrict our attention to type $\mathrm{A}$ and construct a collection of injective homomorphisms among the quantum Grothendieck rings of different ranks. (Of course, by composing them with the isomorphisms between type $\mathrm{B}_{n}$ and type $\mathrm{A}_{2n-1}$ in [6, 7], we can extend our collection to a collection of injective homomorphisms among type $\mathrm{AB}$.) To be more precise, let $\mathscr{C}_{n}$ be a monoidal skeleton of the category $\mathscr{C}$ (= the subcategory $\mathscr{C}_{\mathbb{Z}}$ in the sense of Hernandez–Leclerc [9]) for the quantum loop algebra of $\mathfrak{sl}_{n}$, and $K_{t}(\mathscr{C}_{n})$ its quantum Grothendieck ring. Let $1,n]\coloneqq\{1,2,\ldots,n\}$ be the integer interval between $1$ and $n$, and $I_{n}\coloneqq[1,n-1]$ an index set of simple roots of $\mathfrak{sl}_{n}$. The main result of this paper is the following.
Theorem 1.1 (= Theorem 4.7).
[…]
Combined with the aforementioned result of Nakajima [20], Theorem 1.1 implies that the classical limit of $\Psi_{{\tilde{\xi}},\nu,\xi}$ respects the simple classes. In particular, we find that Brito–Chari’s inflation respects the simple classes, which gives a proof of [1, Conjecture 3.2].
In the last section of this paper, we briefly discuss a categorification of our quantum inflations. By Kang–Kashiwara–Kim [13], the quantum Grothendieck ring $K_{t}(\mathscr{C}_{n})$ is known to be isomorphic to the Grothendieck ring $K(\mathscr{T}_{n})$ of a certain localization $\mathscr{T}_{n}$ of the category of finite-dimensional graded modules over the quiver Hecke algebras of type $\mathrm{A}_{\infty}$. In the last section, we observe the following.
Theorem 1.2 (= Theorem 5.6).
For any choice of $(\xi,{\tilde{\xi}},\nu)$, there is a graded exact monoidal functor $F_{{\tilde{\xi}},\nu,\xi}\colon\mathscr{T}_{n}\to\mathscr{T}_{\tilde{n}}$ categorifying $\Psi_{{\tilde{\xi}},\nu,\xi}$.
Organization
This paper is organized as follows. We recall some fundamental facts on $\mathscr{C}_{n}$ and Brito–Chari’s inflation in §2. We give a brief review on the quantum Grothendieck ring $K_{t}(\mathscr{C}_{n})$ and its presentation in §3. The main part of this paper is §4, where we define the quantum inflations in §4.1, prove their compatibility with the canonical bases in §4.2, and verify Brito–Chari’s conjecture in §4.3. We discuss the categorification of the quantum Grothendieck ring $K_{t}(\mathscr{C}_{n})$ and the quantum inflations in the final §5.
2. Preliminaries
In this section, after fixing our notation in §2.1, we quickly review some basic facts on the representation theory of the quantum loop algebra of type $\mathrm{A}$ in §2.2. Then, we recall the definition of the inflations introduced by Brito–Chari [1] and their conjecture in §2.3.
2.1. Notation
Throughout the paper, we use the following conventions. For a mathematical statement $P$, we set $\delta(P)$ to be $1$ or 0 according that $P$ is true or false. As a special case, we set $\delta_{i,j}\coloneqq\delta(i=j)$. We denote by $\mathbb{Z},\mathbb{Q}$, and $\mathbb{C}$ the sets of integers, rational numbers, and complex numbers respectively. For $a,b\in\mathbb{Z}$ with $a\leq b$, we set $a,b]\coloneqq\{k\in\mathbb{Z}\mid a\leq k\leq b\}$. Given $n\in\mathbb{Z}$, we write $a\equiv_{n}b$ when $a-b\in n\mathbb{Z}$.
Fix $n\in\mathbb{Z}_{>1}$. Let $P_{n}=\bigoplus_{i\in[1,n]}\mathbb{Z}\epsilon_{i}$ be a free abelian group of rank $n$ equipped with a standard bilinear pairing $(\epsilon_{i},\epsilon_{j})=\delta_{i,j}$. The symmetric group $\mathfrak{S}_{n}$ of degree $n$ acts on $P_{n}$ from the left in the natural way. For $a,b\in[1,n]$ with $a<b$, we put $\alpha_{a,b}\coloneqq\epsilon_{a}-\epsilon_{b}\in P_{n}$. The set $R_{n}^{+}\coloneqq\{\alpha_{a,b}\mid a,b\in[1,n],a<b\}$ is identical to the set of positive roots of type $\mathrm{A}_{n-1}$. Let $I_{n}\coloneqq[1,n-1].$ For each $i\in I_{n}$, let $\alpha_{i}\coloneqq\alpha_{i,i+1}\in P_{n}$ be the $i$-th simple root and $s_{i}\in\mathfrak{S}_{n}$ the $i$-th simple reflection (the transposition of $i$ and $i+1$). Let $c_{i,j}\coloneqq(\alpha_{i},\alpha_{j})$ be the Cartan integer for $i,j\in I_{n}$.
We consider the set
$\widehat{I}_{n}\coloneqq\{(i,p)\in I_{n}\times\mathbb{Z}\mid p\equiv_{2}i\}$
and the ring of Laurent polynomials
$\mathcal{Y}_{n}\coloneqq\mathbb{Z}[Y_{i,p}^{\pm 1}\mid(i,p)\in\widehat{I}_{n}].$
Let $\mathcal{M}_{n}^{+}$ (resp. $\mathcal{M}_{n}$) be the subset of $\mathcal{Y}_{n}$ consisting of all the monomials (resp. Laurent monomials) in the variables $Y_{i.p}$, $(i,p)\in\widehat{I}_{n}$. It is an abelian monoid (resp. group) by the multiplication. We refer to an element $m\in\mathcal{M}_{n}^{+}$ as a dominant monomial.
2.2. Quantum loop algebras and $q$-characters
Let $U_{q}(L\mathfrak{sl}_{n})$ be the quantum loop algebra, associated with the complex simple Lie algebra $\mathfrak{sl}_{n}$ of type $\mathrm{A}_{n-1}$. This is a Hopf algebra over $\mathbb{C}$. We refer to [2, Chapter 12] for its precise definition. In this paper, we assume that $q$ is not a root of unity.
Recall that simple finite-dimensional modules of type $\mathbf{1}$ over $U_{q}(L\mathfrak{sl}_{n})$ are parametrized by the multiplicative monoid $(1+z\mathbb{C}[z])^{I_{n}}$ of Drinfeld polynomials up to isomorphism by the fundamental result of Chari–Pressley [2, 3]. For $\pi\in(1+z\mathbb{C}[z])^{I_{n}}$, we denote by $L(\pi)$ the corresponding simple $U_{q}(L\mathfrak{sl}_{n})$-module. We regard the monoid $\mathcal{M}_{n}^{+}$ of dominant monomials as a submonoid of $(1+z\mathbb{C}[z])^{I_{n}}$ through the correspondence
$Y_{i,p}\mapsto(1-\delta_{i,j}q^{p}z)_{j\in I_{n}}.$
Then we consider the category $\mathscr{C}_{n}$ of finite-dimensional $U_{q}(L\mathfrak{sl}_{n})$-modules whose composition factors are isomorphic to $L(m)$’s for $m\in\mathcal{M}_{n}^{+}$. This category is the same as Hernandez–Leclerc’s category $\mathscr{C}_{\mathbb{Z}}$ for $U_{q}(L\mathfrak{sl}_{n})$ in [9]. (Note that the symbol $\mathscr{C}_{n}$ has a different meaning in [9].) The category $\mathscr{C}_{n}$ is closed under taking tensor products and duals. Therefore, it carries a natural structure of rigid monoidal abelian category.
Let $K(\mathscr{C}_{n})$ be the Grothendieck ring of the category $\mathscr{C}_{n}$. It comes with the free $\mathbb{Z}$-basis $\{[L(m)]\mid m\in\mathcal{M}_{n}^{+}\}$ of simple isomorphism classes. The $q$-character map, in the sense of Frenkel–Reshetikhin [5], induces an injective ring homomorphism
$\chi_{q}\colon K(\mathscr{C}_{n})\to\mathcal{Y}_{n}.$
By Frenkel–Mukhin [4], the image of $\chi_{q}$ coincides with the intersection
(2.1) $\bigcap_{i\in I_{n}}\mathbb{Z}[Y_{i,p}(1+A_{i,p+1}^{-1}),Y_{j,s}^{\pm 1}\mid j\neq i,p\equiv_{2}i,s\equiv_{2}j],$
where $A_{i,p}\coloneqq Y_{i,p-1}Y_{i,p+1}Y^{-1}_{i-1,p}Y^{-1}_{i+1,p}$ with $Y_{0,p}=Y_{n,p}=1$ by convention. In what follows, we often identify $K(\mathscr{C}_{n})$ with the image of $\chi_{q}$.
A simple module of the form $L(Y_{i,p})$ for some $(i,p)\in\widehat{I}_{n}$ is called a fundamental module. As a commutative ring, $K(\mathscr{C}_{n})$ is freely generated by the classes of fundamental modules.
2.3. Brito–Chari’s inflation
Let $n,{\tilde{n}}\in\mathbb{Z}_{>1}$ satisfying ${\tilde{n}}/n\in\mathbb{Z}$. In [1], Brito–Chari introduced a injective ring homomorphism
[…]
given by the assignment $\chi_{q}(L(Y_{i,p}))\mapsto\chi_{q}(L(Y_{{\tilde{n}}i/n,{\tilde{n}}p/n}))$ for all $(i,p)\in\widehat{I}_{n}$. The homomorphism $\Psi_{{\tilde{n}},n}$ is called the inflation.
Conjecture 2.1 (Brito–Chari [1, Conjecture 3.2]).
The inflation $\Psi_{{\tilde{n}},n}$ respects the simple isomorphism classes. More precisely, letting
$\psi_{{\tilde{n}},n}\colon\mathcal{M}_{n}^{+}\to\mathcal{M}_{{\tilde{n}}}^{+}$
be the homomorphism given by $Y_{i,p}\mapsto Y_{{\tilde{n}}i/n,{\tilde{n}}p/n}$ for $(i,p)\in\widehat{I}_{n}$, we have
(2.2) $\Psi_{{\tilde{n}},n}([L(m)])=[L(\psi_{{\tilde{n}},n}(m))]$
for any $m\in\mathcal{M}_{n}^{+}$.
In [1, Theorem 2], Brito–Chari verifies the relation (2.2) for some special classes of simple modules $L(m)$ such as (usual and higher order) Kirillov–Reshetikhin modules, snake modules, and arbitrary simple modules in a level-one subcategory $\mathscr{C}_{\xi}$ associated with an increasing height function $\xi$ as in [10].
Later in §4.3, we give a proof of Conjecture 2.1 by constructing an analog of the inflation $\Psi_{{\tilde{n}},n}$ between the quantum Grothendieck rings and showing that it respects their canonical bases.
3. Quantum Grothendieck rings of type $\mathrm{A}$
In this section, we recall the definition of the quantum Grothendieck ring of the category $\mathscr{C}_{n}$ and their canonical bases following Nakajima [20] and Hernandez [8]. We also recall its presentation as the bosonic extension of the half of the quantized enveloping algebra of $\mathfrak{sl}_{n}$ due to Hernandez–Leclerc [11]. The presentation depends on the choice of a height function on the Dynkin diagram of $\mathfrak{sl}_{n}$ and plays a key role in our construction of the quantum analog of inflation in §4.
3.1. Definition
Let $z$ be a formal parameter and $C(z)\coloneqq(\frac{z^{c_{i,j}}-z^{-c_{i,j}}}{z-z^{-1}})_{i,j\in I_{n}}$ a deformed Cartan matrix. For any $i,j\in I_{n}$, the $(i,j)$-entry $\tilde{C}_{i,j}(z)$ of the inverse matrix $C(z)^{-1}$ is a rational function in $z$. We write its Laurent expansion at $z=0$ as
$\tilde{C}_{i,j}(z)=\sum_{k\in\mathbb{Z}}\tilde{c}_{i,j}(k)z^{k}.$
It turns out that all the coefficients $\tilde{c}_{i,j}(k)$ are integers. It is known that we have $\tilde{c}_{i,j}(k)=0$ if $k\leq 0$ or $k\equiv_{2}i+j$ (see [11, §2]).
Define the skew-symmetric bilinear map $\mathscr{N}\colon\mathcal{M}_{n}\times\mathcal{M}_{n}\to\mathbb{Z}$ of abelian groups by $\mathscr{N}(Y_{i,p},Y_{j,s})=\mathscr{N}_{i,j}(p-s)$, where
$\mathscr{N}_{i,j}(k)\coloneqq\tilde{c}_{i,j}(k-1)-\tilde{c}_{i,j}(k+1)-\tilde{c}_{i,j}(-k-1)+\tilde{c}_{i,j}(-k+1),$
for any $i,j\in I_{n}$ and $k\in\mathbb{Z}$. An explicit formula of $\mathscr{N}_{i,j}(k)$ is given in Example 3.8 below.
Let $t$ be another formal parameter with a square root $t^{1/2}$. We endow the free $\mathbb{Z}[t^{\pm 1/2}]$-module
$\mathcal{Y}_{n,t}\coloneqq\mathbb{Z}[t^{\pm 1/2}]\otimes_{\mathbb{Z}}\mathcal{Y}_{n}=\bigoplus_{m\in\mathcal{M}_{n}}\mathbb{Z}[t^{\pm 1/2}]m$
with a $\mathbb{Z}[t^{\pm 1/2}]$-bilinear product * by $m*m^{\prime}\coloneqq t^{\mathscr{N}(m,m^{\prime})/2}mm^{\prime}$ for $m,m^{\prime}\in\mathcal{M}_{n}$. The resulting $\mathbb{Z}[t^{\pm 1/2}]$-algebra $\mathcal{Y}_{n,t}=(\mathcal{Y}_{n,t},*)$ is a quantum torus, which gives a non-commutative deformation of $\mathcal{Y}_{n}$. The specialization at $t^{1/2}=1$ yields a ring homomorphism $\mathop{\mathrm{ev}}\nolimits_{t=1}\colon\mathcal{Y}_{n,t}\to\mathcal{Y}_{n}$. In what follows, we often identify an element $y\in\mathcal{Y}_{n}$ with $1\otimes y\in\mathcal{Y}_{n,t}$. It gives an inclusion $\mathcal{Y}_{n}=1\otimes\mathcal{Y}_{n}\subset\mathcal{Y}_{n,t}$ of a subgroup (not of a subalgebra).
We have the natural anti-involution $y\mapsto\overline{y}$ of $\mathcal{Y}_{n,t}$ given by $\overline{t^{1/2}}=t^{-1/2}$ and $\overline{m}=m$ for all $m\in\mathcal{M}_{n}$. We call it the bar-involution. Each element of the subgroup $\mathcal{Y}_{n}\subset\mathcal{Y}_{n,t}$ is fixed by this bar-involution.
For each $i\in I_{n}$, let $K_{t,i}$ be the $\mathbb{Z}[t^{\pm 1/2}]$-subalgebra of the quantum torus $\mathcal{Y}_{n,t}$ generated by the set
$\{Y_{i,p}(1+A_{i,p+1}^{-1})\mid p\equiv_{2}i\}\cup\{Y_{j,s}^{\pm 1}\mid j\neq i,s\equiv_{2}j\}.$
Following Hernandez [8], we define the quantum Grothendieck ring $K_{t}(\mathscr{C}_{n})$ to be the intersection of them:
$K_{t}(\mathscr{C}_{n})\coloneqq\bigcap_{i\in I_{n}}K_{t,i}.$
Compare to (2.1). By construction, $K_{t}(\mathscr{C}_{n})$ is a $\mathbb{Z}[t^{\pm 1/2}]$-subalgebra of $\mathcal{Y}_{n,t}$. By [20, Corollary 3.6] or [8, Theorem 6.2], we have
$\mathop{\mathrm{ev}}\nolimits_{t=1}(K_{t}(\mathscr{C}_{n}))=\chi_{q}(K(\mathscr{C}_{n}))\simeq K(\mathscr{C}_{n}).$
3.2. Canonical bases
By [12, Proposition 9.3 & Theorem 9.4], the $q$-character of a fundamental module $\chi_{q}(L(Y_{i,p}))$ belongs to $K_{t}(\mathscr{C}_{n})$. This is a special feature of type $\mathrm{A}$ (and $\mathrm{B}$).
Lemma 3.1.
For $(i,p),(j,s)\in\widehat{I}_{n}$, we have
$\chi_{q}(L(Y_{i,p}))*\chi_{q}(L(Y_{j,s}))=\chi_{q}(L(Y_{j,s}))*\chi_{q}(L(Y_{i,p}))\quad\text{if $|p-s|<|i-j|+\delta_{i,j}$.}$
Proof.
This follows from [6, Lemma 9.11(3)]. See also Remark 3.2 below. ∎
Write a dominant monomial $m\in\mathcal{M}_{n}^{+}$ as $m=Y_{i_{1},p_{1}}\cdots Y_{i_{d},p_{d}}$ with $p_{1}\geq\cdots\geq p_{d}$. Then, we set
(3.1) $E_{t}(m)\coloneqq t^{-\sum_{1\leq k<l\leq d}\mathscr{N}(Y_{i_{k},p_{k}},Y_{i_{l},p_{l}})/2}\chi_{q}(L(Y_{i_{1},p_{1}}))*\cdots*\chi_{q}(L(Y_{i_{d},p_{d}})).$
Since $\chi_{q}(L(Y_{i,p}))$ and $\chi_{q}(L(Y_{j,p}))$ mutually commute with respect to * by Lemma 3.1, and $\mathscr{N}_{i,j}(0)=0$ for any $i,j\in I_{n}$, the element $E_{t}(m)$ is well-defined (independent of the ordering of the factors of $m$). The set $\{E_{t}(m)\mid m\in\mathcal{M}_{n}^{+}\}$ gives a free $\mathbb{Z}[t^{\pm 1/2}]$-basis of $K_{t}(\mathscr{C}_{n})$, called the standard basis.
Remark 3.2.
In [8], the standard basis element $E_{t}(m)$ is defined in terms of the element $F_{t}(Y_{i,p})$ such that $Y_{i,p}$ is the unique dominant monomial occurring in $F_{t}(Y_{i,p})$, instead of $\chi_{q}(L(Y_{i,p}))$. Our definition of $E_{t}(m)$ is equivalent to the one in loc. cit. as $F_{t}(Y_{i,p})=\chi_{q}(L(Y_{i,p}))$ holds in type $\mathrm{A}$. Note that the last equality is not true in general.
Theorem 3.3 (Nakajima [20], Hernandez [8], see also [12, Remark 7.8]).
For each dominant monomial $m\in\mathcal{M}_{n}^{+}$, there exists a unique element $\chi_{q,t}(L(m))$ of $K_{t}(\mathscr{C}_{n})$ satisfying
$\overline{\chi_{q,t}(L(m))}=\chi_{q,t}(L(m))\quad\text{and}\quad\chi_{q,t}(L(m))-E_{t}(m)\in\sum_{m^{\prime}\in\mathcal{M}_{n}^{+}}t^{-1}\mathbb{Z}[t^{-1}]E_{t}(m^{\prime}).$
Moreover, the set $\{\chi_{q,t}(L(m))\mid m\in\mathcal{M}_{n}^{+}\}$ forms a free $\mathbb{Z}[t^{\pm 1/2}]$-basis of $K_{t}(\mathscr{C}_{n})$, called the canonical basis of $K_{t}(\mathscr{C}_{n})$.
Remark 3.4.
The transition matrix from the standard basis to the canonical basis turns out to be unitriangular with respect to the so-called Nakajima partial ordering of dominant monomials. In fact, the original defining condition of the canonical basis in [20, 8] requires this unitriangularity as well. The weaker condition in Theorem 3.3 is enough to characterize the basis, as explained in [12, Remark 7.8].
The element $\chi_{q,t}(L(m))$ is called the $t$-analog of $q$-character (or the $(q,t)$-character for short) of the simple module $L(m)$. For a fundamental module $L(Y_{i,p})$, we have $\chi_{q,t}(L(Y_{i,p}))=\chi_{q}(L(Y_{i,p})).$
The following fundamental result was proved by Nakajima using the geometry of quiver varieties.
Theorem 3.5 (Nakajima [20]).
For each $m\in\mathcal{M}_{n}^{+}$, we have
$\mathop{\mathrm{ev}}\nolimits_{t=1}\chi_{q,t}(L(m))=\chi_{q}(L(m)).$
3.3. Presentations
We say that a function $\xi\colon I_{n}\to\mathbb{Z}$ is a height function if it satisfies $\xi(1)\equiv_{2}1$ and $|\xi(i)-\xi(i+1)|=1$ for all $i\in[1,n-2]$. A height function $\xi$ defines a Coxeter element $\tau_{\xi}\in\mathfrak{S}_{n}$ by $\tau_{\xi}\coloneqq s_{i_{1}}s_{i_{2}}\cdots s_{i_{n-1}}$, where we chose a total ordering $I_{n}=\{i_{1},i_{2},\ldots,i_{n-1}\}$ satisfying $\xi(i_{1})\leq\xi(i_{2})\leq\cdots\leq\xi(i_{n-1})$. The element $\tau_{\xi}$ does not depend on the choice of such a total ordering. Following [11, §2], we recursively define the bijection
$\phi_{\xi}\colon\widehat{I}_{n}\xrightarrow{\sim}R_{n}^{+}\times\mathbb{Z}$
by the following two requirements:
  1. (1)
    we have $\phi_{\xi}(i,\xi(i))=(\sum_{j\in I_{n},\xi_{j}-\xi_{i}=|j-i|}\alpha_{j},0)$ for any $i\in I_{n}$,
  2. (2)
    if $\phi_{\xi}(i,p)=(\alpha,k)$, we have
[…]
Remark 3.6.
Our $\tau_{\xi}$ is the same as $\tau^{-1}$ in [11, §2]. Our bijection $\phi_{\xi}$ is the same as the bijection $\varphi$ in loc. cit.
For future use, we recall a relation between the bilinear form $\mathscr{N}$ and the bijection $\phi_{\xi}$.
Lemma 3.7.
Let $\xi\colon I_{n}\to\mathbb{Z}$ be a height function. For any $(i,p),(j,s)\in\widehat{I}_{n}$ with $(i,p)\neq(j,s)$, we have
$\mathscr{N}(Y_{i,p},Y_{j,s})=(-1)^{k+l+\delta(p\geq s)}(\alpha,\beta),$
where $(\alpha,k)=\phi_{\xi}(i,p)$ and $(\beta,l)=\phi_{\xi}(j,s)$. Moreover, we have
$\mathscr{N}(Y_{i,p},Y_{j,s})=0\quad\text{if $p\geq s$ and $k<l$.}$
Proof.
This follows from [11, Proposition 3.2] and [6, Lemma 9.11(2)]. ∎
Example 3.8.
The most important case for the purpose of this paper is when our height function $\xi$ is increasing, namely when it satisfies
(3.2) $\xi(i)=i+2c\quad\text{for any $i\in I_{n}$}$
with a constant $c\in\mathbb{Z}$. In this case, the Coxeter element $\tau_{\xi}$ is $s_{1}s_{2}\cdots s_{n-1}$, which acts on $P_{n}$ by $\epsilon_{i}\mapsto\epsilon_{i+1}$ for all $i\in[1,n]$. Here $\epsilon_{n+1}\coloneqq\epsilon_{1}$ by convention. The bijection $\phi_{\xi}$ is easy to compute as follows. For an integer $p\in\mathbb{Z}$, we write it as
(3.3) $p=k_{n}(p)n+r_{n}(p)\quad\text{with $k_{n}(p)\in\mathbb{Z}$ and $r_{n}(p)\in[1,n]$}.$
Then, we have
$\phi_{\xi}(i,\xi(i)+2p)=\begin{cases}(\alpha_{r_{n}(p),r_{n}(p+i)},2k_{n}(p)+1)&\text{if $r_{n}(p)<r_{n}(p+i)$},\\ (\alpha_{r_{n}(p+i),r_{n}(p)},2k_{n}(p)+2)&\text{if $r_{n}(p)>r_{n}(p+i)$},\end{cases}$
for any $i\in I_{n}$ and $p\in\mathbb{Z}$. Therefore, the inverse bijection $\phi_{\xi}^{-1}$ is given as
(3.4) $\begin{cases}\phi_{\xi}^{-1}(\alpha_{a,b},2k+1)=(b-a,\xi(b-a)+2(kn+a)),\\ \phi_{\xi}^{-1}(\alpha_{a,b},2k+2)=(n+a-b,\xi(n+a-b)+2(kn+b)).\end{cases}$
With these computations, Lemma 3.7 yields an explicit formula of $\mathscr{N}_{i,j}(k)$: for any $i,j\in I_{n}$ and $k\geq\delta_{i,j}$, we have
$\mathscr{N}_{i,j}(k)=\delta(k\equiv_{2n}i+j)-\delta(k\equiv_{2n}i-j)-\delta(k\equiv_{2n}-i+j)+\delta(k\equiv_{2n}-i-j).$
Now, we recall a presentation of the quantum Grothendieck ring $K_{t}(\mathscr{C}_{n})$, which depends on the choice of height function.
Definition 3.9 (Bosonic extension).
Let $\widehat{\mathcal{A}}_{n}$ be a $\mathbb{Q}(t^{1/2})$-algebra presented by the generators $\{e_{i,k}\mid i\in I_{n},k\in\mathbb{Z}\}$ with the following relations:
  1. (R1)
    For any $k\in\mathbb{Z}$ and any $i,j\in I_{n}$, we have
    $\begin{cases}e_{i,k}^{2}e_{j,k}-(t+t^{-1})e_{i,k}e_{j,k}e_{i,k}+e_{j,k}e_{i,k}^{2}=0&\text{if $|i-j|=1$},\\ e_{i,k}e_{j,k}-e_{j,k}e_{i,k}=0&\text{if $|i-j|>1$}.\end{cases}$
  2. (R2)
    For any $k,k^{\prime}\in\mathbb{Z}$ with $k<k^{\prime}$ and $i,j\in I_{n}$, we have
    $e_{i,k}e_{j,k^{\prime}}=t^{(-1)^{k+k^{\prime}}c_{i,j}}e_{j,k^{\prime}}e_{k,i}+(1-t^{-2})\delta_{(i,k),(j,k^{\prime}-1)}.$
The algebra $\widehat{\mathcal{A}}_{n}$ is often called the bosonic extension of the positive half $U_{n}^{+}$ of the quantized enveloping algebra of $\mathfrak{sl}_{n}$ (see §4.1).
Put
$K_{t}(\mathscr{C}_{n})_{loc}\coloneqq K_{t}(\mathscr{C}_{n})\otimes_{\mathbb{Z}[t^{\pm 1/2}]}\mathbb{Q}(t^{1/2}).$
Theorem 3.10 (Hernandez–Leclerc [11], F.–Hernandez–Oh–Oya [6]).
  1. (1)
    Let $\xi\colon I_{n}\to\mathbb{Z}$ be a height function. There is an isomorphism
    $\Psi_{\xi}\colon\widehat{\mathcal{A}}_{n}\xrightarrow{\sim}K_{t}(\mathscr{C}_{n})_{loc}$
    of $\mathbb{Q}(t^{1/2})$-algebras satisfying
    $\Psi_{\xi}(e_{j,k})=\chi_{q,t}(L(Y_{i,p}))\quad\text{if $\phi_{\xi}(i,p)=(\alpha_{j},k)$}$
    for any $(j,k)\in I_{n}\times\mathbb{Z}$.
  2. (2)
    Let $\xi^{\prime}\colon I_{n}\to\mathbb{Z}$ be another height function. The automorphism $\Psi_{\xi,\xi^{\prime}}\coloneqq\Psi_{\xi}\circ\Psi_{\xi^{\prime}}^{-1}$ of $K_{t}(\mathscr{C}_{n})_{loc}$ induces a permutation of the simple $(q,t)$-characters. In other words, there is a certain permutation $\psi_{\xi,\xi^{\prime}}$ of the set $\mathcal{M}_{n}^{+}$ such that we have
    $\Psi_{\xi,\xi^{\prime}}(\chi_{q,t}(L(m)))=\chi_{q,t}(L(\psi_{\xi,\xi^{\prime}}(m)))$
    for any $m\in\mathcal{M}_{n}^{+}$. In particular, $\Psi_{\xi,\xi^{\prime}}$ restricts to a $\mathbb{Z}[t^{\pm 1/2}]$-algebra automorphism of $K_{t}(\mathscr{C}_{n})$.
Remark 3.11.
In general, the permutation $\psi_{\xi,\xi^{\prime}}$ of $\mathcal{M}_{n}^{+}$ in Theorem 3.10 (2) is only piecewise linear, not a monoid automorphism. By construction, it satisfies $\psi_{\xi,\xi^{\prime}}(Y_{i,p})=Y_{j,s}$ if $\phi_{\xi^{\prime}}(i,p)=\phi_{\xi}(j,s)=(\alpha,k)$ with $\alpha$ being a simple root. But it is not always true if $\alpha$ is a non-simple root.
Remark 3.12.
If $\xi-\xi^{\prime}=2c$ for some constant $c\in\mathbb{Z}$, the automorphism $\Psi_{\xi,\xi^{\prime}}$ coincides with the spectral parameter shift $Y_{i,p}\mapsto Y_{i,p+2c}$. In general, $\Psi_{\xi,\xi^{\prime}}$ can be written as a composition of the braid symmetries $\sigma_{i}$, $i\in I_{n}$, introduced in [14], which we use in the proof of Proposition 4.3 below, and a spectral parameter shift. See [7, the paragraph before Proposition 6.2].
4. Quantum inflations
[…]
4.1. Construction
Let $n\in\mathbb{Z}_{>1}$ and $U_{n}^{+}$ the positive half of the quantized enveloping algebra of $\mathfrak{sl}_{n}$ over $\mathbb{Q}(t^{1/2})$. Namely, $U_{n}^{+}$ is the $\mathbb{Q}(t^{1/2})$-algebra presented by the generators $\{e_{i}\mid i\in I_{n}\}$ with the quantum Serre relations:
$\begin{cases}e_{i}^{2}e_{j}-(t+t^{-1})e_{i}e_{j}e_{i}+e_{j}e_{i}^{2}=0&\text{if $|i-j|=1$},\\ e_{i}e_{j}-e_{j}e_{i}=0&\text{if $|i-j|>1$}.\end{cases}$
For each $k\in\mathbb{Z}$, the assignment $e_{i}\mapsto e_{i,k}$ gives rise to a $\mathbb{Q}(t^{1/2})$-algebra homomorphism
$\iota_{k}\colon U_{n}^{+}\to\widehat{\mathcal{A}}_{n}.$
In the sequel, we use the following $t$-commutator notation:
$x,y]_{t^{1/2}}\coloneqq t^{1/2}xy-t^{-1/2}yx.$
For any $a,b\in[1,n]$ with $a<b$, we define the element $e(\alpha_{a,b})$ of $U_{n}^{+}$ by
(4.1) $e(\alpha_{a,b})\coloneqq\frac{[e_{a},[e_{a+1},\cdots[e_{b-2},e_{b-1}]_{t^{1/2}}\cdots]_{t^{1/2}}]_{t^{1/2}}}{(t-t^{-1})^{b-a-1}}.$
We have $e(\alpha_{i})=e_{i}$ for any $i\in I_{n}$ by definition. Note that $e(\alpha_{a,b})$ and $e(\alpha_{c,d})$ commute if $b<c$.
Lemma 4.1.
For any $n\in\mathbb{Z}_{>1}$ and $a,b,c\in[1,n]$ with $a<b<c$, we have
$e(\alpha_{a,c})=\frac{[e(\alpha_{a,b}),e(\alpha_{b,c})]_{t^{1/2}}}{t-t^{-1}}.$
Proof.
For $x,y,z\in U_{n}^{+}$, we have $x,y]_{t^{1/2}},z]_{t^{1/2}}=[x,[y,z]_{t^{1/2}}]_{t^{1/2}}$ if $xz=zx$. With this remark in mind, the assertion follows by induction on cac-a. ∎
Let $n,{\tilde{n}}\in\mathbb{Z}_{>1}$ with $n<{\tilde{n}}$, and $\nu\colon[1,n]\to[1,{\tilde{n}}$] a (strictly) increasing function. Consider an assignment from $\{e_{i,k}\mid i\in I_{n},k\in\mathbb{Z}\}\subset\widehat{\mathcal{A}}_{n}$ to $\widehat{\mathcal{A}}_{\tilde{n}}$ given by
(4.2) $e_{i,k}\mapsto\iota_{k}(e(\alpha_{\nu(i),\nu(i+1)}))\in\widehat{\mathcal{A}}_{\tilde{n}}\quad\text{for each $i\in I_{n}$ and $k\in\mathbb{Z}$}.$
Example 4.2.
For example, when $n=2$ and ${\tilde{n}}=3$, there are three increasing functions $\nu\colon[1,2]\to[1,3]$. The assignment (4.2) is written as
$e_{1,k}\mapsto\begin{cases}e_{1,k}&\text{if $(\nu(1),\nu(2))=(1,2)$},\\ (t-t^{-1})^{-1}[e_{1,k},e_{2,k}]_{t^{1/2}}&\text{if $(\nu(1),\nu(2))=(1,3)$},\\ e_{2,k}&\text{if $(\nu(1),\nu(2))=(2,3)$},\end{cases}$
for each $k\in\mathbb{Z}$.
Proposition 4.3.
For any increasing function $\nu\colon[1,n]\to[1,{\tilde{n}}$], the assignment (4.2) gives rise to a $\mathbb{Q}(t^{1/2})$-algebra homomorphism
$\Upsilon_{\nu}\colon\widehat{\mathcal{A}}_{n}\to\widehat{\mathcal{A}}_{\tilde{n}}$.
Proof.
For the sake of simplicity, we put ${\tilde{e}}_{i,k}\coloneqq\iota_{k}(e(\alpha_{\nu(i),\nu(i+1)}))$ in this proof. We have to verify that the collection $\{{\tilde{e}}_{i,k}\mid i\in I_{n},k\in\mathbb{Z}\}\subset\widehat{\mathcal{A}}_{\tilde{n}}$ satisfies the defining relations (R1) & (R2) of the algebra $\widehat{\mathcal{A}}_{n}$. We proceed by induction on ${\tilde{n}}-n\in\mathbb{Z}_{>0}$.
First, we consider the case when ${\tilde{n}}=n+1$. Note that, for any increasing function $\nu\colon[1,n]\to[1,n+1]$, there is a unique $l\in[0,n]$ such that we have $\nu(i)=i+\delta(i>l)$ for all $i\in I_{n}$. Then, we have
(4.3) ${\tilde{e}}_{i,k}=\begin{cases}e_{i,k}&\text{if $i<l$},\\ (t-t^{-1})^{-1}[e_{l,k},e_{l+1,k}]_{t^{1/2}}&\text{if $i=l$},\\ e_{i+1,k}&\text{if $i>l$}.\end{cases}$
The desired relations (R1) & (R2) for ${\tilde{e}}_{i,k}$ and ${\tilde{e}}_{j,k^{\prime}}$ with $i,j\in I_{n}\setminus\{l\}$ follows immediately from this formula. For the remaining relations, we can check them by using the braid group symmetry generated by a collection $\{\sigma_{l}^{\pm 1}\mid l\in I_{n+1}\}$ of $\mathbb{Q}(t^{1/2})$-algebra automorphisms of $\widehat{\mathcal{A}}_{n+1}$ introduced by Kashiwara–Kim–Oh–Park in [14]. Recall that these automorphisms are given by the formula
$\sigma_{l}(e_{i,k})$ $=\begin{cases}(t-t^{-1})^{-1}[e_{i,k},e_{l,k}]_{t^{1/2}}&\text{if $|i-l|=1$},\\ e_{i,k+\delta_{i,l}}&\text{otherwise},\end{cases}$
$\sigma^{-1}_{l}(e_{i,k})$ $=\begin{cases}(t-t^{-1})^{-1}[e_{l,k},e_{i,k}]_{t^{1/2}}&\text{if $|i-l|=1$},\\ e_{i,k-\delta_{i,l}}&\text{otherwise}.\end{cases}$
See [14, Theorem 2.3] (also [16, Theorem 3.1] for more general statement with a detailed proof). Comparing the formula with (4.3), we have
${\tilde{e}}_{i,k}=\begin{cases}\sigma_{l+1}(e_{i,k})&\text{if $i\leq l$},\\ \sigma_{l}^{-1}(e_{i+1,k})&\text{if $i\geq l$}.\end{cases}$
Using this, the relations (R1) & (R2) for ${\tilde{e}}_{i,k}$ and ${\tilde{e}}_{j,k^{\prime}}$ with $i=l$ or $j=l$ can be checked immediately. Thus, $\Upsilon_{\nu}$ is well-defined when ${\tilde{n}}=n+1$.
For the induction step, we note that an increasing function $\nu\colon[1,n]\to[1,{\tilde{n}}$] with ${\tilde{n}}-n>1$ can be factorized as $\nu=\nu^{\prime\prime}\circ\nu^{\prime}$, where $\nu^{\prime}\colon[1,n]\to[1,n^{\prime}]$ and $\nu^{\prime\prime}\colon[1,n^{\prime}]\to[1,{\tilde{n}}$] are some increasing functions with $n^{\prime}\in[n+1,{\tilde{n}}-1]$. By induction hypothesis, the $\mathbb{Q}(t^{1/2})$-algebra homomorphisms $\Upsilon_{\nu^{\prime}}$ and $\Upsilon_{\nu^{\prime\prime}}$ are well-defined. We apply Lemma 4.1 repeatedly to find that
${\tilde{e}}_{i,k}=(\Upsilon_{\nu^{\prime\prime}}\circ\Upsilon_{\nu^{\prime}})(e_{i,k})$
holds for any $i\in I_{n}$ and $k\in\mathbb{Z}$. This implies that the collection $\{{\tilde{e}}_{i,k}\mid i\in I_{n},k\in\mathbb{Z}\}$ satisfies the desired relations (R1) & (R2). Thus, the homomorphism $\Upsilon_{\nu}$ is also well-defined and $\Upsilon_{\nu}=\Upsilon_{\nu^{\prime\prime}}\circ\Upsilon_{\nu^{\prime}}$ holds. ∎
Definition 4.4 (Quantum inflations).
Let $n,{\tilde{n}}\in\mathbb{Z}_{>1}$. For any height functions $\xi\colon I_{n}\to\mathbb{Z},{\tilde{\xi}}\colon I_{\tilde{n}}\to\mathbb{Z}$ and any increasing function $\nu\colon[1,n]\to[1,{\tilde{n}}$], we define the quantum inflation
$\Psi_{{\tilde{\xi}},\nu,\xi}\colon K_{t}(\mathscr{C}_{n})_{loc}\to K_{t}(\mathscr{C}_{\tilde{n}})_{loc}$
to be the composition $\Psi_{{\tilde{\xi}},\nu,\xi}\coloneqq\Psi_{{\tilde{\xi}}}\circ\Upsilon_{\nu}\circ\Psi_{\xi}^{-1}$, where $\Psi_{\xi}$ and $\Psi_{{\tilde{\xi}}}$ are the isomorphisms from Theorem 3.10.
Remark 4.5.
  1. (1)
    When $n={\tilde{n}}$, the homomorphism $\Upsilon_{\nu}$ is the identity automorphism of $K_{t}(\mathscr{C}_{n})$ and we have $\Psi_{{\tilde{\xi}},\nu,\xi}=\Psi_{{\tilde{\xi}},\xi}$ in the notation of Theorem 3.10 (2).
  2. (2)
    Let $n_{1},n_{2},n_{3}\in\mathbb{Z}_{>1}$ with $n_{1}\leq n_{2}\leq n_{3}$. For any height functions $\xi_{i}\colon I_{n_{i}}\to\mathbb{Z}$, $i\in\{1,2,3\}$, and increasing functions $\nu_{i}\colon[1,n_{i}]\to[1,n_{i+1}]$, $i\in\{1,2\}$, we have $\Psi_{\xi_{3},\nu_{2},\xi_{2}}\circ\Psi_{\xi_{2},\nu_{1},\xi_{1}}=\Psi_{\xi_{3},\nu_{2}\circ\nu_{1},\xi_{1}}.$ This follows from the construction and the equality $\Upsilon_{\nu_{2}}\circ\Upsilon_{\nu_{1}}=\Upsilon_{\nu_{2}\circ\nu_{1}}$ as we saw in the proof of Proposition 4.3.
Remark 4.6.
When ${\tilde{n}}=n+1$ with $\nu(1)=1$ and $\nu(n)={\tilde{n}}$, the homomorphism $\Upsilon_{\nu}\colon\widehat{\mathcal{A}}_{n}\to\widehat{\mathcal{A}}_{n+1}$ can be thought of the bosonic extension of an edge contraction $U_{n}^{+}\to U_{n+1}^{+}$ in the sense of Li [19]. In general, $\Upsilon_{\nu}$ is the bosonic extension of an iteration of edge contractions.
4.2. Compatibility with canonical bases
We prove that the quantum inflations respect the canonical bases of the quantum Grothendieck rings. We retain the notation from the previous subsection.
Theorem 4.7.
[…]
(4.4) $\Psi_{{\tilde{\xi}},\nu.\xi}(\chi_{q,t}(L(m)))=\chi_{q,t}(L(\psi_{{\tilde{\xi}},\nu,\xi}(m)))$
for any $m\in\mathcal{M}_{n}^{+}$. In particular, $\Psi_{{\tilde{\xi}},\nu,\xi}$ restricts to an injective $\mathbb{Z}[t^{\pm 1/2}]$-algebra homomorphism
[…]
compatible with the bar-involutions.
Proof.
Let $\xi^{\prime}\colon I_{n}\to\mathbb{Z}$ and ${\tilde{\xi}}^{\prime}\colon I_{\tilde{n}}\to\mathbb{Z}$ be another pair of height functions. By Remark 4.5, we have $\Psi_{{\tilde{\xi}},\nu,\xi}=\Psi_{{\tilde{\xi}},{\tilde{\xi}}^{\prime}}\circ\Psi_{{\tilde{\xi}}^{\prime},\nu,\xi^{\prime}}\circ\Psi_{\xi^{\prime},\xi}$. Therefore, as soon as we prove the statement for $\Psi_{{\tilde{\xi}}^{\prime},\nu,\xi^{\prime}}$ with an injective map $\psi_{{\tilde{\xi}}^{\prime},\nu,\xi^{\prime}}\colon\mathcal{M}_{n}^{+}\to\mathcal{M}_{\tilde{n}}$, the statement for $\Psi_{{\tilde{\xi}},\nu,\xi}$ also follows with the map $\psi_{{\tilde{\xi}},\nu,\xi}\coloneqq\psi_{{\tilde{\xi}},{\tilde{\xi}}^{\prime}}\circ\psi_{{\tilde{\xi}}^{\prime},\nu,\xi^{\prime}}\circ\psi_{\xi^{\prime},\xi}$, thanks to Theorem 3.10 (2). Thus, it suffices to consider the special case when both height functions $\xi$ and $\tilde{\xi}$} are increasing, which is treated in Proposition 4.8 below. ∎
With a given increasing function $\nu\colon[1,n]\to[1,{\tilde{n}}$], we associate an injective map $\nu_{*}\colon R_{n}^{+}\to R_{\tilde{n}}^{+}$ by
$\nu_{*}(\alpha_{a,b})\coloneqq\alpha_{\nu(a),\nu(b)}\quad\text{for any $a,b\in[1,n]$ with $a<b$}.$
Proposition 4.8.
Assume that both $\xi\colon I_{n}\to\mathbb{Z}$ and ${\tilde{\xi}}\colon I_{\tilde{n}}\to\mathbb{Z}$ are increasing height functions. Then, the equality (4.4) holds for any $m\in\mathcal{M}_{n}^{+}$ with $\psi_{{\tilde{\xi}},\nu,\xi}\colon\mathcal{M}_{n}^{+}\to\mathcal{M}_{\tilde{n}}^{+}$ being the injective monoid homomorphism given by
$\psi_{{\tilde{\xi}},\nu,\xi}(Y_{i,p})\coloneqq Y_{j,s}\quad\text{if $(j,s)=(\phi_{\tilde{\xi}}^{-1}\circ(\nu_{*}\times\mathsf{id}_{\mathbb{Z}})\circ\phi_{\xi})(i,p)$}$
for each $(i,p)\in\widehat{I}_{n}$.
Proof.
Recall that a height function $\xi\colon I_{n}\to\mathbb{Z}$ defines a Dynkin quiver $Q_{\xi}$ of type $\mathrm{A}_{n-1}$ so that we have an arrow $i\to j$ in $Q_{\xi}$ if and only if |ij|=1|i-j|=1 and $\xi(i)>\xi(j)$. For example, when $\xi$ is increasing, $Q_{\xi}$ is a monotone quiver: $Q_{\xi}=(1\leftarrow 2\leftarrow\cdots\leftarrow n-1)$. Associated to a Dynkin quiver $Q$ of type $\mathrm{A}_{n-1}$, the rescaled dual PBW generator $\tilde{E}^{*}_{Q}(\alpha)\in U_{n}^{+}$ is defined for each positive root $\alpha\in R_{n}^{+}$ as in [11, §6.1]. (Note that $\tilde{E}^{*}_{Q}(\alpha)$ is written simply as $\tilde{E}^{*}(\alpha)$ in loc. cit. and our $t$ is equal to $v$ in loc. cit.) For an arbitrary height function $\xi$, by the construction of the isomorphism $\Psi_{\xi}$ and [11, Theorem 6.1], we know that the isomorphism $\Psi_{\xi}\colon\widehat{\mathcal{A}}_{n}\to K_{t}(\mathscr{C}_{n})_{loc}$ sends $\iota_{k}(\tilde{E}^{*}_{Q_{\xi}}(\alpha))$ to $\chi_{q,t}(L(Y_{i,p}))$ with $(i,p)=\phi_{\xi}^{-1}(\alpha,k)$ for any $(\alpha,k)\in R_{n}^{+}\times\mathbb{Z}$.
Now, we assume that $\xi$ is increasing. Then, it is straightforward to check that the rescaled PBW generator $\tilde{E}_{Q_{\xi}}^{*}(\alpha)$ coincides with the element $e(\alpha)$ defined in (4.1) for any $\alpha\in R_{n}^{+}$. Therefore, in this case, we have $\Psi_{\xi}(\iota_{k}(e(\alpha)))=\chi_{q,t}(L(Y_{i,p}))$ with $(i,p)=\phi_{\xi}^{-1}(\alpha,k)$ for any $(\alpha,k)\in R_{n}^{+}\times\mathbb{Z}$.
Thus, assuming that both $\xi$ and $\tilde{\xi}$} are increasing, we have
(4.5) $\Psi_{{\tilde{\xi}},\nu,\xi}(\chi_{q,t}(L(Y_{i,p})))=\chi_{q,t}(L(\psi_{{\tilde{\xi}},\nu,\xi}(Y_{i,p})))$
for any $(i,p)\in\widehat{I}_{n}$. In addition, since $(\alpha,\beta)=(\nu_{*}\alpha,\nu_{*}\beta)$ holds for any $\alpha,\beta\in R_{n}^{+}$, Lemma 3.7 yields
(4.6) $\mathscr{N}(Y_{i,p},Y_{j,s})=\mathscr{N}(\psi_{{\tilde{\xi}},\nu,\xi}(Y_{i,p}),\psi_{{\tilde{\xi}},\nu,\xi}(Y_{j,s}))$
for any $(i,p),(j,s)\in\widehat{I}_{n}$. From (4.5) and (4.6), together with Lemma 3.1, it follows that
$\Psi_{{\tilde{\xi}},\nu,\xi}(E_{t}(m))=E_{t}(\psi_{{\tilde{\xi}},\nu,\xi}(m))$
for all $m\in\mathcal{M}_{n}^{+}$ (recall the definition of $E_{t}(m)$ in (3.1)). Moreover, it is immediate to see from the construction that $\Psi_{{\tilde{\xi}},\nu,\xi}$ commutes with the bar-involutions. Therefore, for any $m\in\mathcal{M}_{n}^{+}$, we find that the element $\Psi_{{\tilde{\xi}},\nu,\xi}(\chi_{q,t}(L(m)))$ satisfies the two characterizing properties of $\chi_{q,t}(L(\psi_{{\tilde{\xi}},\nu,\xi}(m)))$ in Theorem 3.3. Thus, we obtain the desired equality (4.4). ∎
Remark 4.9.
For general height functions $\xi$ and $\tilde{\xi}$}, the injective map $\psi_{{\tilde{\xi}},\nu,\xi}\colon\mathcal{M}_{n}^{+}\to\mathcal{M}_{\tilde{n}}^{+}$ is not a monoid homomorphism. It is only piecewise linear. See Remark 3.11.
[…]
(4.7) ${K_{t}(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}},\nu,\xi}}$$\scriptstyle{\mathop{\mathrm{ev}}\nolimits_{t=1}}$${K_{t}(\mathscr{C}_{\tilde{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathop{\mathrm{ev}}\nolimits_{t=1}}$${K(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}},\nu,\xi}|_{t=1}}$${K(\mathscr{C}_{\tilde{n}}).}$
Corollary 4.10.
For any choice of $(\xi,{\tilde{\xi}},\nu)$, the homomorphism $\Psi_{{\tilde{\xi}},\nu,\xi}|_{t=1}$ induces an injective map from the set of simple classes of $\mathscr{C}_{n}$ to that of $\mathscr{C}_{\tilde{n}}$. More precisely, we have
$\Psi_{{\tilde{\xi}},\nu,\xi}|_{t=1}([L(m)])=[L(\psi_{{\tilde{\xi}},\nu,\xi}(m))]$
for any $m\in\mathcal{M}_{n}^{+}$, where $\psi_{{\tilde{\xi}},\nu,\xi}\colon\mathcal{M}_{n}^{+}\to\mathcal{M}_{\tilde{n}}^{+}$ is as in Theorem 4.7.
Proof.
The assertion follows from the commutativity of the diagram (4.7) together with Theorem 3.5 and Theorem 4.7. ∎
4.3. A proof of Conjecture 2.1
Let $n,{\tilde{n}}\in\mathbb{Z}_{>1}$ satisfying ${\tilde{n}}/n\in\mathbb{Z}$ as in §2.3. We define the function $\nu_{{\tilde{n}}/n}\colon[1,n]\to[1,{\tilde{n}}$] by $\nu_{{\tilde{n}}/n}(i)\coloneqq{\tilde{n}}i/n$. We also take the increasing height functions $\xi_{0}\colon I_{n}\to\mathbb{Z}$ and ${\tilde{\xi}}_{0}\colon I_{\tilde{n}}\to\mathbb{Z}$ given by $\xi_{0}(i)\coloneqq i$ and ${\tilde{\xi}}_{0}(i)\coloneqq i$ respectively.
[…]
Proposition 4.11.
We have $\Psi_{{\tilde{\xi}}_{0},\nu_{{\tilde{n}}/n},\xi_{0}}|_{t=1}=\Psi_{{\tilde{n}},n}$ and $\psi_{{\tilde{\xi}}_{0},\nu_{{\tilde{n}}/n},\xi_{0}}=\psi_{{\tilde{n}},n}$.
Proof.
A direct computation using the formula (3.4) in Example 3.8 shows that the homomorphism $\psi_{{\tilde{\xi}}_{0},\nu_{{\tilde{n}}/n},\xi_{0}}$ sends $Y_{i,p}$ to $Y_{{\tilde{n}}i/n,{\tilde{n}}p/n}$ for all $(i,p)\in\widehat{I}_{n}$. Hence, we get the equality $\psi_{{\tilde{\xi}}_{0},\nu_{{\tilde{n}}/n},\xi_{0}}=\psi_{{\tilde{n}},n}$. Then, by Corollary 4.10, we have $\Psi_{{\tilde{\xi}}_{0},\nu_{{\tilde{n}}/n},\xi_{0}}|_{t=1}([L(Y_{i,p})])=[L(Y_{{\tilde{n}}i/n,{\tilde{n}}p/n})]$ for any $(i,p)\in\widehat{I}_{n}$, which proves the equality $\Psi_{{\tilde{\xi}}_{0},\nu_{{\tilde{n}}/n},\xi_{0}}|_{t=1}=\Psi_{{\tilde{n}},n}$. ∎
5. Categorification
We end this paper with a brief discussion on a categorification of our quantum inflations using the quiver Hecke algebras of type $\mathrm{A}_{\infty}$. It is built on the works of Kang–Kashiwara–Kim [13], Kashiwara–Park [17], and Kashiwara–Kim–Oh–Park [14].
5.1. Quiver Hecke algebras of type $\mathrm{A}_{\infty}$
Let $P\coloneqq\bigoplus_{a\in\mathbb{Z}}\mathbb{Z}\epsilon_{a}$ be a free abelian group of countably infinite rank equipped with the standard bilinear pairing $(\epsilon_{a},\epsilon_{b})=\delta_{a,b}$. For each $a,b\in\mathbb{Z}$ with $a<b$, we set $\alpha_{a,b}\coloneqq\epsilon_{a}-\epsilon_{b}$ and $\alpha_{a}\coloneqq\alpha_{a,a+1}$. The set $R^{+}\coloneqq\{\alpha_{a,b}\mid a,b\in\mathbb{Z},a<b\}$ is identical to the set of positive roots of type $\mathrm{A}_{\infty}$. Let $Q^{+}\coloneqq\sum_{a\in\mathbb{Z}}\mathbb{Z}_{\geq 0}\alpha_{a}\subset P$.
For each $\beta\in Q^{+}$, the quiver Hecke algebra $R(\beta)$ of type $\mathrm{A}_{\infty}$ is defined. This is an associative $\mathbb{Z}$-graded algebra over $\mathbb{C}$ given by an explicit presentation. See [13]. Let $\mathscr{A}_{\beta}$ denote the category of finite-dimensional $\mathbb{Z}$-graded $R(\beta)$-modules and set $\mathscr{A}\coloneqq\bigoplus_{\beta\in Q^{+}}\mathscr{A}_{\beta}$. For $M\in\mathscr{A}_{\beta}$ and $M^{\prime}\in\mathscr{A}_{\beta}^{\prime}$, we can form their convolution product $M\circ M^{\prime}\in\mathscr{A}_{\beta+\beta^{\prime}}$. This gives rise to a bi-exact functor $\mathscr{A}_{\beta}\times\mathscr{A}_{\beta^{\prime}}\to\mathscr{A}_{\beta+\beta^{\prime}}$, which makes the category $\mathscr{A}$ into a $\mathbb{Z}$-graded monoidal abelian category. In particular, the Grothendieck ring $K(\mathscr{A})\coloneqq\bigoplus_{\beta\in Q^{+}}K(\mathscr{A}_{\beta})$ is a $\mathbb{Z}[t^{\pm 1}]$-algebra, where $t^{\pm 1}$ are the grading shifts. By convention, our $t$ is the same as $q^{-1}$ in [13]. Taking the graded dual of the modules gives an anti-involution $M\mapsto M^{*}$ of $\mathscr{A}$. It satisfies
(5.1) $(t^{\pm 1}M)^{*}\simeq t^{\mp 1}M^{*}\quad\text{and}\quad(M\circ M^{\prime})^{*}\simeq t^{-(\beta,\beta^{\prime})}M^{\prime*}\circ M^{*}$
for $M\in\mathscr{A}_{\beta}$ and $M^{\prime}\in\mathscr{A}_{\beta^{\prime}}$. The Grothendieck ring $K(\mathscr{A})$ has a free $\mathbb{Z}[t^{\pm 1}]$-basis formed by the classes of self-dual simple modules. Any simple module in $\mathscr{A}$ is isomorphic to a grading shift of a self-dual one.
For each positive root $\alpha\in R^{+}$, there is a self-dual one-dimensional module $L(\alpha)\in\mathscr{A}(\alpha)$. In [13], $L(\alpha_{a,b})$ was denoted by L(a,b1)L(a,b-1). The set $\{[L(\alpha)]\mid\alpha\in R^{+}\}$ generates the $\mathbb{Z}[t^{\pm 1/2}]$-algebra $K(\mathscr{A})$ and satisfies the relation
(5.2) $(1-t^{-2})[L(\alpha_{a,b})]=[L(\alpha_{a,b})\circ L(\alpha_{b,c})]-t^{-1}[L(\alpha_{b,c})\circ L(\alpha_{a,b})]$
for any $a,b,c\in\mathbb{Z}$ with $a<b<c$. See [13, Proposition 4.3(vi)].
5.2. Categorification of $K_{t}(\mathscr{C}_{n})$
Given $n\in\mathbb{Z}_{>1}$, let $S_{n}$ de the automorphism of $P$ given by $S_{n}(\epsilon_{a})=\epsilon_{a+n}$ for $a\in\mathbb{Z}$. Define a bilinear form $B_{n}(\cdot,\cdot)\colon P\times P\to\mathbb{Z}$ by
$B_{n}(x,y)\coloneqq\sum_{k\in\mathbb{Z}_{>0}}(S_{n}^{k}(x),y).$
Then, twisting the convolution product $\circ$ in $\mathscr{A}$, we set
(5.3) $M\star M^{\prime}\coloneqq t^{B_{n}(\beta,\beta^{\prime})}M\circ M^{\prime}$
for $M\in\mathscr{A}_{\beta}$ and $M^{\prime}\in\mathscr{A}_{\beta^{\prime}}$. This defines another monoidal structure on $\mathscr{A}$, and we write $\mathscr{A}_{n}\coloneqq(\mathscr{A},\star)$ for the $\mathbb{Z}$-graded monoidal abelian category thus obtained. From (5.1) and (5.3), we have
(5.4) $(M\star M^{\prime})^{*}\simeq t^{-(r_{n}(\beta),r_{n}(\beta^{\prime}))}M^{\prime*}\star M^{*},$
for $M\in\mathscr{A}_{\beta}$ and $M^{\prime}\in\mathscr{A}_{\beta^{\prime}}$, where $r_{n}\colon P\to P_{n}=\bigoplus_{i\in[1,n]}\mathbb{Z}\epsilon_{i}$ is the homomorphism given by $r_{n}(\epsilon_{a})\coloneqq\epsilon_{r_{n}(a)}$ in the notation of (3.3).
In [13, §4], a certain $\mathbb{Z}$-graded rigid monoidal abelian category $\mathscr{T}_{n}$ was constructed as a localization of the category $\mathscr{A}_{n}$. By definition, it comes with an exact monoidal functor $\Omega_{n}\colon\mathscr{A}_{n}\to\mathscr{T}_{n}$ satisfying the following properties and it is universal among such monoidal functors:
  1. (1)
    $\Omega_{n}(L(\alpha_{a,b}))=0$ if ba>nb-a>n,
  2. (2)
    for any $a\in\mathbb{Z}$, there is an isomorphism between $\Omega_{n}(L(\alpha_{a,a+n}))$ and the unit object of $\mathscr{T}_{n}$ satisfying a certain compatibility condition (see [13, Proposition A.12] for detail).
Moreover, $\Omega_{n}$ sends simple modules in $\mathscr{A}_{n}$ to simple objects or zeros of $\mathscr{T}_{n}$, and every simple object of $\mathscr{T}_{n}$ is the image of a simple module in $\mathscr{A}_{n}$ under $\Omega_{n}$. See [13, Proposition A.11]. We put $L_{n}(\alpha_{a,b})\coloneqq\Omega_{n}(L(\alpha_{a,b}))\neq 0$ for $a,b\in\mathbb{Z}$ with $b-a\in I_{n}=[1,n-1]$. The monoidal category $\mathscr{T}_{n}$ gives a categorification of the quantum Grothendieck ring $K_{t}(\mathscr{C}_{n})$ as follows.
Theorem 5.1 (Kang–Kashiwara–Kim [13]).
There is an isomorphism
$\Phi_{n}\colon K(\mathscr{T}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\xrightarrow{\sim}K_{t}(\mathscr{C}_{n})$
of $\mathbb{Z}[t^{\pm 1/2}]$-algebras satisfying
(5.5) $\Phi_{n}(t^{1/2}[L_{n}(\alpha_{a,b})])=\chi_{q,t}(L(Y_{b-a,a+b}))$
for any $a,b\in\mathbb{Z}$ with $b-a\in I_{n}$. Moreover, there exists an exact monoidal functor $F_{n}\colon\mathscr{T}_{n}\to\mathscr{C}_{n}$ which sends simple objects of $\mathscr{T}_{n}$ to simple objects of $\mathscr{C}_{n}$ and satisfies $\mathop{\mathrm{ev}}\nolimits_{t=1}\circ\Phi_{n}=[F_{n}]$.
Proof.
This follows from [13, Theorem 4.32] and its proof. The functor $F_{n}$ arises from the quantum affine Schur–Weyl duality. ∎
Although the following corollary may be well-known for the experts, we shall give a detailed proof as the author could not find a proper reference.
Corollary 5.2.
The isomorphism $\Phi_{n}$ in Theorem 5.1 sends each simple class in $K(\mathscr{T}_{n})$ to a $t^{\frac{1}{2}\mathbb{Z}}$-multiple of a simple $(q,t)$-character in $K_{t}(\mathscr{C}_{n})$.
Proof.
As our base field $\mathbb{C}$ is of characteristic zero, the basis of the self-dual simple modules of $K(\mathscr{A})$ correspond to the dual canonical basis of the quantum unipotent coordinate ring of type $\mathrm{A}_{\infty}$ by [21, 23]. In particular, the self-dual simple classes are parametrized by the set $(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$ (equivalent to the set of Zelevinsky’s multi-segments) as follows. Define a total ordering $\leq$ of the set $R^{+}$ so that we have $\alpha_{a,b}<\alpha_{a^{\prime},b^{\prime}}$ if $a<a^{\prime}$, or $a=a^{\prime}$ & $b<b^{\prime}$. For each $\mathbf{c}=(c_{\alpha})_{\alpha\in R^{+}}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$, we define
(5.6) $M(\mathbf{c})=t^{-\sum_{\alpha\in R^{+}}c_{\alpha}(c_{\alpha}-1)/2}\mathop{\circ}_{\alpha\in R^{+}}^{\leftarrow}L(\alpha)^{\circ c_{\alpha}},$
[…]
(5.7) $L(\mathbf{c})]-[M(\mathbf{c})]\in\sum_{\mathrm{wt}(\mathbf{c}^{\prime})=\mathrm{wt}(\mathbf{c})}t^{-1}\mathbb{Z}[t^{-1}][M(\mathbf{c}^{\prime})],$
where $\mathrm{wt}(\mathbf{c})\coloneqq\sum_{\alpha\in R^{+}}c_{\alpha}\alpha\in Q^{+}$. See [18, Proposition 16].
Let $\sigma$ be the involution of $K(\mathscr{A}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]$ given by $\sigma(t^{1/2})=t^{-1/2}$ and $\sigma([M])=t^{(r_{n}(\beta),r_{n}(\beta))/2}[M^{*}]$ for $M\in\mathscr{A}_{\beta}$. The equation (5.4) tells us that $\sigma$ is an anti-involution of $K(\mathscr{A}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]$. For $M\in\mathscr{A}_{\beta}$, we define its rescaled class $\widetilde{M}]\coloneqq t^{(r_{n}(\beta),r_{n}(\beta))/4}[M]$ so that $\widetilde{M}$] is fixed by $\sigma$ if and only if $M]=[M^{*}]$. We have
$\widetilde{L(\alpha_{a,b})}]=t^{(1-\delta_{b-a,n})/2}[L(\alpha_{a,b})]$
for any $a,b\in\mathbb{Z}$ with $b-a\in[1,n]$. Letting
$\Phi_{n}^{\prime}\coloneqq\Phi_{n}\circ([\Omega_{n}]\otimes 1)\colon K(\mathscr{A}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\to K_{t}(\mathscr{C}_{n}),$
[…]
Observe that, if $\alpha_{a,b}\geq\alpha_{a^{\prime},b^{\prime}}$ and $b-a,b^{\prime}-a^{\prime}\in[1,n]$, we have
$B_{n}(\alpha_{a,b},\alpha_{a^{\prime},b^{\prime}})=-\delta(a=a^{\prime}=b-n=b^{\prime}-n),$
and hence
$\widetilde{M(\mathbf{c})}]=t^{\sum_{\alpha>\alpha^{\prime}}c_{\alpha}c_{\alpha^{\prime}}(r_{n}(\alpha),r_{n}(\alpha^{\prime}))/2}\mathop{\star}^{\leftarrow}_{\alpha\in R^{+}}[\widetilde{L(\alpha)}]^{\star c_{\alpha}}$
for any $\mathbf{c}\in(\mathbb{Z}_{\geq 0})^{\oplus R^{+}}$ such that $c_{\alpha_{a,b}}=0$ if ba>nb-a>n. We also note that
$(r_{n}(\alpha_{a,b}),r_{n}(\alpha_{a^{\prime},b^{\prime}}))=-\mathscr{N}(Y_{b-a,a+b},Y_{b^{\prime}-a^{\prime},a^{\prime}+b^{\prime}})$
holds if $\alpha_{a,b}>\alpha_{a^{\prime},b^{\prime}}$ and $b-a,b^{\prime}-a^{\prime}\in I_{n}$ by Lemma 3.7 and the computation in Example 3.8. Therefore, by Theorem 5.1 and Lemma 3.1, we have
$\Phi^{\prime}_{n}([\widetilde{M(\mathbf{c})}])=\begin{cases}0&\text{if $c_{\alpha_{a,b}}\neq 0$ for some $a,b\in\mathbb{Z}$ with $b-a>n$},\\ E_{t}(m_{\mathbf{c}})&\text{otherwise}\end{cases}$
[…]
Remark 5.3.
As far as the author understands, such a categorification of the quantum Grothendieck ring is known only for type $\mathrm{A}$ at this moment. One of the special features of type $\mathrm{A}$ is that the bozonic extension algebra $\widehat{\mathcal{A}}_{n}$ can be obtained as an explicit quotient of the half of the quantized enveloping algebra of type $\mathrm{A}_{\infty}$, which enables us to construct the categorification out of the quiver Hecke algebras of type $\mathrm{A}_{\infty}$. It would be very interesting if we have a similar categorification for the other types.
5.3. Categorification of quantum inflations
Fix $n,{\tilde{n}}\in\mathbb{Z}_{>1}$ with $n<{\tilde{n}}$. Let $\nu\colon[1,n]\to[1,{\tilde{n}}$] be an increasing function as before. We extend $\nu$ to an increasing function $\nu\colon\mathbb{Z}\to\mathbb{Z}$ by
$\nu(p)\coloneqq k_{n}(p){\tilde{n}}+\nu(r_{n}(p))$
for any $p\in\mathbb{Z}$, where we used the notation from (3.3). Then, we have the induced homomorphism $\nu_{*}\colon P\to P$ given by $\nu_{*}(\epsilon_{a})\coloneqq\epsilon_{\nu(a)}$, which restricts to $R^{+}\to R^{+}$ satisfying $\nu_{*}(\alpha_{a,b})=\alpha_{\nu(a),\nu(b)}$ for any $a,b\in\mathbb{Z}$ with $a<b$. Let us consider the collection $\{L(\nu_{*}(\alpha_{a}))\mid a\in\mathbb{Z}\}$ of simple modules in $\mathscr{A}$, and apply the construction of Kashiwara–Park [17] to the duality datum formed by the affinizations of these simple modules. Note that the degree of the normalized $R$-matrix between $L(\nu_{*}(\alpha_{a}))$ and $L(\nu_{*}(\alpha_{b}))$ is equal to $(\nu_{*}(\alpha_{a}),\nu_{*}(\alpha_{b}))=-(\alpha_{a},\alpha_{b})$ as computed in [13, Proposition 4.3.(vii)]. The duality datum gives rise to a $\mathbb{Z}$-graded exact monoidal functor
$\widehat{F}_{\nu}\colon\mathscr{A}\to\mathscr{A}$
satisfying $\widehat{F}_{\nu}(L(\alpha_{a}))\simeq L(\nu_{*}(\alpha_{a}))$ for any $a\in\mathbb{Z}$. By (5.2), it follows that $\widehat{F}_{\nu}(L(\alpha))\simeq L(\nu_{*}(\alpha))$ for any $\alpha\in R^{+}$. Since $B_{{\tilde{n}}}(\nu_{*}(x),\nu_{*}(y))=B_{n}(x,y)$ for any $x,y\in P$, the functor $\widehat{F}_{\nu}$ intertwines the twisted convolution products, and hence gives a $\mathbb{Z}$-graded exact monoidal functor
$\widehat{F}_{\nu}\colon\mathscr{A}_{n}\to\mathscr{A}_{\tilde{n}}$.
By the universality of the localizations (see the paragraph before Theorem 5.1), there exists a unique $\mathbb{Z}$-graded exact monoidal functor
$F_{\nu}\colon\mathscr{T}_{n}\to\mathscr{T}_{\tilde{n}}$
such that $F_{\nu}\circ\Omega_{n}\simeq\Omega_{\tilde{n}}\circ\widehat{F}_{\nu}$. It satisfies $F_{\nu}(L_{n}(\alpha_{a,b}))\simeq L_{\tilde{n}}(\alpha_{\nu(a),\nu(b)})$ for any $a,b\in\mathbb{Z}$ with $b-a\in I_{n}$.
Proposition 5.4.
Let $\xi_{0}\colon I_{n}\to\mathbb{Z}$ and ${\tilde{\xi}}_{0}\colon I_{\tilde{n}}\to\mathbb{Z}$ be the height functions given by $\xi_{0}(i)=i$ and ${\tilde{\xi}}_{0}(i)=i$ as in §4.3. For any increasing function $\nu\colon[1,n]\to[1,{\tilde{n}}$], the following diagram commutes:
(5.8) ${K(\mathscr{T}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{n}}\scriptstyle{\sim}$$\scriptstyle{[F_{\nu}]\otimes 1}$${K_{t}(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}}_{0},\nu,\xi_{0}}}$${K(\mathscr{T}_{\tilde{n}})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\tilde{n}}}\scriptstyle{\sim}$${K_{t}(\mathscr{C}_{\tilde{n}}).}$
In particular, $F_{\nu}$ sends simple objects of $\mathscr{T}_{n}$ to simple objects of $\mathscr{T}_{\tilde{n}}$.
Proof.
In view of Theorem 5.1, it is enough to prove that $\Psi_{{\tilde{\xi}}_{0},\nu,\xi_{0}}$ sends $\chi_{q,t}(L(Y_{b-a,a+b}))$ to $\chi_{q,t}(L(Y_{\nu(b)-\nu(a),\nu(a)+\nu(b)}))$ for each $a,b\in\mathbb{Z}$ with $b-a\in I_{n}$. This is easily checked by Proposition 4.8 and the explicit computation of $\phi_{\xi_{0}}$ and $\phi_{{\tilde{\xi}}_{0}}$ in Example 3.8. ∎
Remark 5.5.
Specializing (5.8) at $t=1$, we get the commutative diagram:
${K(\mathscr{T}_{n})/(t-1)K(\mathscr{T}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[F_{n}]}\scriptstyle{\sim}$$\scriptstyle{[F_{\nu}]}$${K(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}}_{0},\nu,\xi_{0}}|_{t=1}}$${K(\mathscr{T}_{\tilde{n}})/(t-1)K(\mathscr{T}_{\tilde{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[F_{\tilde{n}}]}\scriptstyle{\sim}$${K(\mathscr{C}_{\tilde{n}}).}$
The composition $F_{\tilde{n}}\circ F_{\nu}\colon\mathscr{T}_{n}\to\mathscr{C}_{\tilde{n}}$ can be constructed more directly following the recipe of [15, §6.1] based on the generalized quantum affine Schur–Weyl duality in the sense of [13]. In [15, Theorem 6.9], it is also proved that the resulting functor always respects the simple classes. In this way, one may obtain an alternative proof of Conjecture 2.1 without using the quantum Grothendieck rings.
Thus, the homomorphism $\Psi_{{\tilde{\xi}}_{0},\nu,\xi_{0}}$ has a categorical lift $F_{\nu}$. We also have a categorical lift of the homomorphism $\Psi_{{\tilde{\xi}},\nu,\xi}$ for general height functions $\xi$ and $\tilde{\xi}$}. To see this, we recall the obvious factorization $\Psi_{{\tilde{\xi}},\nu,\xi}=\Psi_{{\tilde{\xi}},{\tilde{\xi}}_{0}}\circ\Psi_{{\tilde{\xi}}_{0},\nu,\xi_{0}}\circ\Psi_{\xi_{0},\xi}$. As we already mentioned in Remark 3.12, the automorphism $\Psi_{\xi_{0},\xi}$ can be written as a composition of braid symmetries $\sigma_{i}$, $i\in I_{n}$, and a spectral parameter shift $\Sigma_{c}$ for some $c\in\mathbb{Z}$ mapping each $\chi_{q,t}(Y_{i,p})$ to $\chi_{q,t}(L(Y_{i,p+2c}))$. By Kashiwara–Kim–Oh–Park [14, Theorem 3.3], we know that the braid symmetry $\sigma_{i}$ has a categorical lift. As for the spectral parameter shift $\Sigma_{c}$, we have an obvious categorical lift, that is, the auto-equivalence of $\mathscr{T}_{n}$ transforming $L_{n}(\alpha_{a,b})$ to $L_{n}(\alpha_{a+c,b+c})$ for any $a,b\in\mathbb{Z}$ with $b-a\in I_{n}$. Thus, by composition, the automorphism $\Psi_{\xi_{0},\xi}$ has a categorical lift. The same holds for $\Psi_{{\tilde{\xi}},{\tilde{\xi}}_{0}}$ as well. Thus, together with Proposition 5.4, we obtain the following conclusion.
Theorem 5.6.
For any height functions $\xi\colon I_{n}\to\mathbb{Z}$, ${\tilde{\xi}}\colon I_{\tilde{n}}\to\mathbb{Z}$ and any increasing function $\nu\colon[1,n]\to[1,{\tilde{n}}$], there exists a $\mathbb{Z}$-graded exact monoidal functor
$F_{{\tilde{\xi}},\nu,\xi}\colon\mathscr{T}_{n}\to\mathscr{T}_{\tilde{n}}$
making the following diagram commute:
${K(\mathscr{T}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{n}}\scriptstyle{\sim}$$\scriptstyle{[F_{{\tilde{\xi}},\nu,{\tilde{\xi}}}]\otimes 1}$${K_{t}(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}},\nu,\xi}}$${K(\mathscr{T}_{\tilde{n}})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\tilde{n}}}\scriptstyle{\sim}$${K_{t}(\mathscr{C}_{\tilde{n}}).}$
In particular, $F_{{\tilde{\xi}},\nu,\xi}$ sends simple objects of $\mathscr{T}_{n}$ to simple objects of $\mathscr{T}_{\tilde{n}}$.
Acknowledgments
The author would like to thank the anonymous referees for their useful comments and suggestions. This work was supported by JSPS KAKENHI Grant No. JP23K12955.
Declarations
The author has no relevant financial or non-financial interests to disclose. We do not analyze or generate any datasets, because our work proceeds within a mathematical approach.